OpenStudy (anonymous):

1. In 1 kilogram of pure water, the mass of hydrogen atoms is approximately 112 grams, or 0.112 kilograms. What is the energy equivalent of this mass in joules?

OpenStudy (frostbite):

You remember the famous equation made by Einstein for mass–energy equivalence?

OpenStudy (anonymous):

yes @Frostbite

OpenStudy (frostbite):

Perfect! that is the equation we are going to use here.

OpenStudy (anonymous):

E=mc^2 @Frostbite

OpenStudy (frostbite):

Perfect. Now try set into the equation. \[c=299.792,458 ~ \frac{ m }{ s }\]

OpenStudy (frostbite):

(use the kilo unit) as joule is defined as: \[\large J=\frac{ kg \times m ^{2} }{ s ^{2} }\]

OpenStudy (anonymous):

what is the m/8

OpenStudy (frostbite):

m is the mass.

OpenStudy (anonymous):

Ok i couldnt tell if that was an 8 or an s I knew what the M stood for

OpenStudy (anonymous):

ok now what?

OpenStudy (frostbite):

Oh sorry. btw I wrote c wrong (american coma and EU comma :P) \[\large c=299,792,458 ~ \frac{ m }{ s }\] So set into the equation: \[\large E=0.112 \times c ^{2}\] and get E

OpenStudy (anonymous):

I didnt get a correct answer @Frostbite

OpenStudy (frostbite):

What did you get?

OpenStudy (anonymous):

a crazy long number

OpenStudy (anonymous):

@Frostbite

OpenStudy (frostbite):

\[E=0.122 ~ kg \times (299,792,458)^{2} ~ \frac{ m ^{2} }{ s ^{2} }\] You can see it is going to be big.

OpenStudy (frostbite):

Try write what you get.

OpenStudy (anonymous):

1.0964813180589e+16 @Frostbite

OpenStudy (frostbite):

I get the same \(1.01 \times10^{16} ~ J \)

OpenStudy (anonymous):

so that is the answer? @Frostbite

OpenStudy (frostbite):

Use 3 significant digits (best when they only give us 3 significant digits too)

OpenStudy (frostbite):

Yep. It should be.

OpenStudy (anonymous):

Thanks so much. @Frostbite

OpenStudy (frostbite):

No problem at all. Hope it was helpful and similar questions in the future is just right on.